OpenStudy (anonymous):

Find the other endpoint. Endpoint:(-1, 9), midpoint: (-9, -10)

OpenStudy (anonymous):

from -1 to -9 is left 8, 8 more to the left is -17 from 9 to -10 is 19 down, 19 more is -29 your point should be (-17,-29)

OpenStudy (anonymous):

i mean using the midpoint formula though

OpenStudy (anonymous):

then solve \[\frac{a-1}{2}=-9, \frac{b+9}{2}=-10\] for a and b